1-5. Two point charges, q1 = 8 µC and q2 = -5 µC, are separated by a distance r = 0.1 m. What is the magnitude of the electric force? Note: 1 µC = 10-6 C.


1
Expert's answer
2022-03-22T15:48:00-0400

Answer

Electric force is given by

"F=\frac{kq_1q_2}{r}\\=\frac{9*10^9*8*5*10^{-12}}{0.1*0.1}\\=36N"